Your CEO, Will Dewey red about a two risky asset concept that a guy named Hany came up with. He wants you to determine some information on his idea for a two-stock portfolio called ne Boats. (note the standard deviations are given for both firms in the two previous problems). Find the expected mean retum AND the expected standard deviation of the nu Boats portfolio percentage of the portfolio in Rho Boats bwteen rho boats 70.00% and nu news 0.07000 nu mean return 0.12000stdevp ofnu 0.80000 mean retu ofrho Boats Rb 0.08000 std deviation rho Boats Ơrbo
Answer #1

Expected mean return=70%*0.07+30%*0.08=0.073

Standard deviation=sqrt((70%*0.12)^2+(30%*0.20)^2+2*70%*30%*0.12*0.20*0.8)=0.136821051